Ammonia is an important compound and useful to industries such as agricultural, pharmaceuticals, textile, plastics, petroleum, mining, and commercial. In the ammonia production, Haber process is utilized and it is shown as, N2(g) + 3H2(g) ➡️ 2NH3(g).

How many grams of hydrogen gas are needed to produce 10,000.0 g of ammonia?
1
Expert's answer
2020-12-01T09:16:58-0500

Mr(NH3) = Ar(N) + 3*Ar(H) = 14 g/mol + 3*1 g/mol = 17 g/mol

n(NH3) = m(NH3)/Mr(NH3) = (10 g)/(17 g/mol) = 0.5882 mol

n(H2) = n(NH3)*3/2 (stoichiometry coefficients)

n(H2) = (0.5882 mol)*3/2 = 0.8823 mol

m(H2) = n(H2)*Mr(H2) = 2*n(H2)*Ar(H) = 2*(0.8823 mol)*(1 g/mol) = 1.7646 g